A slot machine is configured so that there is a 1/2000 probability of winning the jackpot on any individual trial. Suppose a guest claims that she played the slot machine 5 times and hit the jackpot twice. Find the probability of exactly 2 jackpots in 5 trials. find the probability of at least 2 jackpots in 5 trials. does the guest's claim of hitting 2 jackpots in 5 trials seem valid?
Answer by edjones(8007)

p=1/2000, q=1999/2000, n=5
nCr=5C2=10 coefficient
10p^2*q^3
=.0000025
.
Ed